### v1

# A Critical Study of Automatic Evaluation in Sign Language Translation

Abstract
Automatic evaluation metrics are crucial for advancing sign language translation (SLT). Current SLT evaluation metrics, such as BLEU and ROUGE, are only text-based, and it remains unclear to what extent text-based metrics can reliably capture the quality of SLT outputs. To address this gap, we investigate the limitations of text-based SLT evaluation metrics by analyzing six metrics, including BLEU, chrF, and ROUGE, as well as BLEURT on the one hand, and large language model (LLM)-based evaluators such as G-Eval and GEMBA zero-shot direct assessment on the other hand. Specifically, we assess the consistency and robustness of these metrics under three controlled conditions: paraphrasing, hallucinations in model outputs, and variations in sentence length. Our analysis highlights the limitations of lexical overlap metrics and demonstrates that while LLM-based evaluators better capture semantic equivalence often missed by conventional metrics, they can also exhibit bias toward LLM-paraphrased translations. Moreover, although all metrics are able to detect hallucinations, BLEU tends to be overly sensitive, whereas BLEURT and LLM-based evaluators are comparatively lenient toward subtle cases. This motivates the need for multimodal evaluation frameworks that extend beyond text-based metrics to enable a more holistic assessment of SLT outputs.
Keywords: sign language translation, automatic evaluation, LLM evaluators
[…]
A Critical Study of Automatic Evaluation in Sign Language Translation
Shakib Yazdani1, Yasser Hamidullah1, Cristina España-Bonet1,2,
Eleftherios Avramidis1, Josef van Genabith1
1German Research Center for Artificial Intelligence (DFKI GmbH),
Saarland Informatics Campus, Saarbrücken, Germany
2Barcelona Supercomputing Center (BSC-CNS), Barcelona, Catalonia, Spain
{shakib.yazdani,yasser.hamidullah,cristinae,eleftherios.avramidis,josef.van_genabith}@dfki.de
Abstract content
1.  Introduction
Sign languages are the primary communication systems for millions of deaf and hard-of-hearing people across the globe. They are highly expressive visual languages that convey meaning through hand signs, facial expressions, mouthings, and body posture (Stokoe, 1980). Automatic sign language translation (SLT) aims to translate sign language videos into spoken language text to help bridge the communication gap between the deaf and hearing communities (Camgoz et al., 2018). However, evaluating SLT systems remains a significant challenge, since current evaluation practices rely solely on textual outputs, offering little insight into how effectively the visual-linguistic nature of sign language is captured. Human evaluation, though often the most reliable and serving as the gold standard in evaluating SLT, is costly and time-consuming, making it difficult to scale. As a result, current SLT models typically rely on text-based automatic evaluation metrics. Consequently, a valid measure of translation quality that accounts for this cross-modal grounding is still missing in SLT (Müller et al., 2023). In addition to the limitations of current evaluation metrics, SLT systems are also prone to hallucinations (Zhang et al., 2023; Hamidullah et al., 2025a), fluent but incorrect translations that misrepresent the signed input. Evaluating how metrics respond to such cases is crucial for understanding their reliability.
When translating from sign language to spoken language text, traditional lexical overlap machine translation (MT) metrics such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) are commonly adopted. However, these text-based metrics fall short in capturing the multimodal complexity of sign language. To address this gap, recent studies have proposed evaluation metrics tailored to sign language tasks. SignBLEU (Kim et al., 2024), for instance, operates at the gloss level and aligns better with human judgments, while SiLVERScore (Imai et al., 2025) introduces an embedding-based approach for assessing sign language generation. Yet, SignBLEU depends on gloss annotations, and SiLVERScore has not been evaluated in the context of SLT. Beyond these, embedding-based and large language model (LLM)-based metrics developed for text-to-text MT—such as BLEURT (Sellam et al., 2020) and GEMBA (Kocmi and Federmann, 2023) have shown strong correlations with human judgments in text-to-text MT (Freitag et al., 2024). Nevertheless, these approaches remain underexplored in SLT, and there is still no comprehensive analysis of their strengths and limitations for evaluating SLT systems.
In this work, we systematically examine this evaluation gap through a comparative analysis of three lexical overlap metrics (BLEU, ROUGE, and chrF), one embedding-based metric (BLEURT), and two LLM-based evaluators (G-Eval and GEMBA), applied to four recent off-the-shelf SLT models, one gloss-based (TwoStream-SLT) and three gloss-free (SEM-SLT, SpaMo, and Signformer). The evaluation is performed under three controlled scenarios: paraphrasing (including word reordering), hallucinations in model outputs, and variations in sentence length. We summarize our contributions as follows: (1) we show that automatic lexical overlap metrics such as BLEU, chrF, and ROUGE are sensitive to surface-level lexical variation (paraphrasing) rather than true semantic equivalence. In contrast, embedding-based metric BLEURT and LLM-based evaluators G-Eval and GEMBA better reflect semantic similarity and meaning preservation in SLT outputs; (2) our analysis shows that all metrics are able to reliably distinguish hallucinated from non-hallucinated outputs. However, BLEU is highly sensitive in cases of extreme hallucination, whereas BLEURT and LLM-based evaluators are able to better capture meaning, though they tend to under-penalize subtle hallucinations when the translation remains fluent; (3) a fine-grained analysis by sentence length indicates systematic inconsistencies between metric types: lexical overlap metrics often rank models differently than BLEURT or LLM-based evaluators, highlighting that evaluation outcomes can vary depending on the chosen metric and sentence length.
2.  Related Work
2.1.  Evaluation Metrics
SLT models have traditionally relied on lexical overlap MT metrics for automatic translation. Specifically, BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) have been widely used following the neural SLT work by Camgoz et al. (2018). Following Müller et al. (2022), studies in SLT have adopted the embedding-based metric BLEURT (Sellam et al., 2020). However, these metrics do not take the visual source input into account. Recently, Kim et al. (2024) proposed SignBLEU, an evaluation metric designed to reduce information loss when assessing SLT outputs. By operating at the gloss level, it achieves stronger alignment with human judgments than prior text-based metrics. Nevertheless, SignBLEU remains constrained to the text modality and relies on manually curated gloss annotations, which limits scalability. More recently, Imai et al. (2025) introduced SiLVERScore, an embedding-based metric that evaluates sign language generation in a multimodal semantic space. While it shows promising results, SiLVERScore was not evaluated on SLT.
The progress and advancements in LLMs have motivated research into their potential applications for automated assessment across various tasks. Liu et al. (2023) pioneered LLM-based automated assessment by proposing G-Eval, a framework that uses Chain-of-Thought reasoning (Wei et al., 2022) along with task descriptions and evaluation metrics to evaluate performance on tasks including text summarization and dialogue generation. Following this direction, Kocmi and Federmann (2023) proposed GEMBA, a GPT-based evaluation metric to assess the MT quality both with- and without reference in a zero-shot prompting fashion. Recently, inspired by G-Eval, Tong et al. (2025) extended the framework to image and video caption evaluation using GPT-4o, achieving a high correlation with human judgments.
$2.$  Sign Language Translation
SLT aims to convert sign language videos into spoken language text. SLT models are often categorized into gloss-free and methods with gloss supervision. Camgoz et al. (2018) pioneered SLT by framing it as a neural machine translation (NMT) task, and further extending it by leveraging transformers for end-to-end SLT (Camgoz et al., 2020). MSKA-SLT (Guan et al., 2025) sets a new state-of-the-art on the Phoenix-2014T (Camgoz et al., 2018) and CSL-Daily (Zhou et al., 2021) benchmarks as a gloss-based model, with BLEU scores of $29.03$ and $25.52$, respectively. The approach models keypoints extracted from the face, body, and both hands using an attention mechanism, and is jointly optimized with a CTC loss and a self-distillation loss. However, glosses require manual human annotation, which is labor-intensive and difficult to scale. In contrast, recent approaches attempt to model SLT at the sentence level and a gloss-free fashion. Building on this objective, recent works have incorporated LLMs into SLT frameworks (Gong et al., 2024; Wong et al., 2024; Chen et al., 2024), employed contrastive learning objectives (Cheng et al., 2023), replaced gloss supervision with sentence-level embeddings (Hamidullah et al., 2024), and extended SLT to multilingual scenarios (Yin et al., 2022; Yazdani et al., 2025; Tan et al., 2025; Hamidullah et al., 2025b).
2.3.  Hallucination
A model is said to hallucinate when it generates content that either lacks logical coherence or diverges from the source material, asserting information or events that are unfounded or inconsistent with the known facts (Ji et al., 2023). While hallucination detection has received increasing attention in the MT community, with studies on multilingual comparisons (Guerreiro et al., 2023) and LLM-based approaches for both low- and high-resource languages (Benkirane et al., 2024), research on hallucination in SLT remains limited. Zhang et al. (2023) reported, based on manual analysis, that SLT models such as SLTUNET suffer greatly from hallucinations, where the generated translations often show limited correlation with the sign video. More recently, Hamidullah et al. (2025a) proposed a hallucination detection framework and further demonstrated that gloss-free SLT models are particularly prone to severe hallucinations compared to their gloss-based counterparts. These findings highlight the need for a deeper understanding of how evaluation metrics respond to hallucinated outputs and how robust current SLT models are under such conditions.
3.  Experimental Setup
3.1.  Task and Datasets
In our experiments, we evaluate the limitations of current SLT metrics with respect to paraphrasing (including word reordering), hallucination sensitivity, and variation in translation length. We use the Phoenix-2014T  [^1] Phoenix-2014T was selected because it is the most widely used benchmark in SLT research and contains gloss-level annotations necessary for evaluating gloss-based models. dataset (Camgoz et al., 2018), covering weather forecasts in German sign language (DGS) to evaluate and compare the performance of various models in the above aspects.
Impact of paraphrasing on SLT evaluation.
For this aspect of our study, we compare recent off-the-shelf SLT models (Sec. 3.3) across both traditional evaluation metrics and LLM-based ones (Sec. 3.2) when we paraphrase the translation predicted by the model. We focus on paraphrasing, which naturally captures surface-level variations such as synonym substitution. We use GPT-4o-mini with the following prompt to generate paraphrased versions of model translations, which are then compared against their original references.
Paraphrase the following sentence into anatural and fluent form. Do not alter anynumbers written in words into digits orvice versa — keep the format as it is inthe original text.
For lexical overlap metrics (BLEU, chrF, and ROUGE) and embedding-based BLEURT, and to ensure robustness against paraphrasing variations, we include a multi-reference evaluation in which the model’s translated and paraphrased output is compared against an augmented set of eleven references comprising the original human reference and ten gold-standard paraphrases.
Impact of hallucination on SLT evaluation.
SLT models are prone to hallucinations, where the generated text diverges from the visual input. Yet, there has been no systematic investigation into how such hallucinations affect translation quality or the reliability of automatic SLT evaluation metrics. Motivated by findings from Benkirane et al. (2024), which demonstrate the effectiveness of LLMs for hallucination detection in MT, we employ Llama-3-70B to identify hallucinated outputs in a reference-based setting. [^2] The prompt used for hallucination detection is available in Appendix A.1. To analyze how evaluation metrics behave under varying degrees of hallucination, we consider a Severity Ranking scheme that categorizes outputs into four levels: No Hallucination, Small Hallucination, Partial Hallucination, and Full Hallucination.
Impact of sentence length on SLT evaluation.
In SLT, most studies report only the average evaluation score over the entire dataset, which can obscure important performance variations. To investigate the sensitivity of evaluation metrics to sentence length, we conduct a fine-grained analysis by segmenting the test set based on the number of words per sentence. Specifically, we group sentences into five ranges: 1–6 (42 sentences), 7–12 (286 sentences), 13–18 (220 sentences), 19–24 (78 sentences), and 25–31 (16 sentences). This segmentation allows us to investigate how translation quality varies with sentence complexity and length, providing deeper insights into where existing models struggle and how hallucination patterns may correlate with sentence structure.
Method Setting BLEU chrF BLEURT ROUGE
Signformer Original $14.8\scriptstyle\pm 1.4$ $35.0\scriptstyle\pm 1.2$ $0.425\scriptstyle\pm 0.016$ $32.8\scriptstyle\pm 2.0$
Paraphrased $5.7\scriptstyle\pm 0.8$ $31.7\scriptstyle\pm 0.9$ $0.452\scriptstyle\pm 0.015$ $27.7\scriptstyle\pm 1.8$
Multi-Ref $18.6\scriptstyle\pm 1.4$ $38.8\scriptstyle\pm 1.3$ $0.509\scriptstyle\pm 0.016$ $36.5\scriptstyle\pm 1.8$
SEM-SLT Original $23.7\scriptstyle\pm 2.1$ $45.8\scriptstyle\pm 1.6$ $0.484\scriptstyle\pm 0.012$ $47.9\scriptstyle\pm 1.9$
Paraphrased $10.0\scriptstyle\pm 1.3$ $38.3\scriptstyle\pm 1.1$ $0.537\scriptstyle\pm 0.014$ $38.0\scriptstyle\pm 1.6$
Multi-Ref $26.8\scriptstyle\pm 2.4$ $47.0\scriptstyle\pm 1.7$ $0.594\scriptstyle\pm 0.013$ $49.8\scriptstyle\pm 1.8$
SpaMo Original $22.2\scriptstyle\pm 2.0$ $44.2\scriptstyle\pm 1.2$ $0.542\scriptstyle\pm 0.015$ $43.2\scriptstyle\pm 2.1$
Paraphrased $9.3\scriptstyle\pm 1.0$ $38.6\scriptstyle\pm 0.9$ $0.558\scriptstyle\pm 0.013$ $35.3\scriptstyle\pm 1.7$
Multi-Ref $26.1\scriptstyle\pm 1.8$ $48.1\scriptstyle\pm 1.1$ $0.618\scriptstyle\pm 0.015$ $46.7\scriptstyle\pm 1.3$
TwoStream-SLT Original $28.2\scriptstyle\pm 2.1$ $50.5\scriptstyle\pm 1.4$ $0.597\scriptstyle\pm 0.013$ $50.3\scriptstyle\pm 2.1$
Paraphrased $12.7\scriptstyle\pm 0.8$ $43.3\scriptstyle\pm 0.9$ $0.604\scriptstyle\pm 0.011$ $40.5\scriptstyle\pm 1.6$
Multi-Ref $34.2\scriptstyle\pm 1.6$ $54.5\scriptstyle\pm 1.1$ $0.670\scriptstyle\pm 0.009$ $53.7\scriptstyle\pm 1.4$
Table 1: Performance comparison of SLT models on the Phoenix-2014T test set under original, paraphrased, and multi-reference conditions across multiple evaluation metrics.
3.2.  Evaluation Metrics
For our analysis, we consider the metrics as per Müller et al. (2022, 2023), specifically BLEU [^3] BLEU|nrefs:1|bs:25|tok:none|eff:no| case:mixed|smooth:exp|version:2.5.1 (via SacreBLEU; Post, 2018) for lexical overlap, ROUGE (Lin, 2004) [^4] ROUGE|metrics:rougeL|nrefs:1| stemmer:true|bs:25|version:0.1.2 for recall-oriented n-gram overlap, chrF (Popović, 2015)555chrF|nrefs:1|bs:25|case:mixed|eff:yes| nc:6|nw:2|space:no|version:2.5.1 that uses character $n$-grams, and embedding-based BLEURT (Sellam et al., 2020) [^6] BLEURT v0.0.2 using checkpoint BLEURT-20. for semantic quality. We use bootstrap resampling for statistical significance reporting.
We also extend the traditional range of SLT evaluation metrics by presenting a sign language adapted version of G-Eval (Liu et al., 2023). G-Eval introduces a framework that leverages prompting techniques to generate evaluation scores that strongly correlate with human preferences. The prompt is organized into three modules: (1) evaluation dimensions, which specify the aspects to be judged; (2) step-by-step reasoning, where Chain-of-Thought (CoT) (Wei et al., 2022) guides the LLM through the evaluation procedure; and (3) scoring with references, which constrains the output format and incorporates human translations as ground truth. Following Sato et al. (2024), we prompt the LLMs to evaluate the translation quality of the models in Section 3.3 for Adequacy and Fluency, each rated on a 5-point Likert scale (1155). Following Fomicheva et al. (2022); Kocmi and Federmann (2023), we further include GEMBA zero-shot Direct Assessment (Kocmi and Federmann, 2023), where the LLM is prompted to evaluate each translation hypothesis and provide a quality score ranging from 0 (completely incorrect) to $100$ (perfect translation). Our LLM experimental setup includes four models: GPT-4.1-nano, Qwen3-8B, Llama-3.1-8B, and QWQ-32B. These were chosen to include both a proprietary LLM and also the open-source ones. [^7] All prompts used for G-Eval and GEMBA are provided in Appendices A.2 and A.3.
$3.$  SLT Models
SLT models are either gloss-based or gloss-free. We include three gloss-free models: SpaMo (Hwang et al., 2025), SEM-SLT (Hamidullah et al., 2024), and Signformer (Yang, 2024). We also include TwoStream-SLT (Chen et al., 2022b) as a gloss-based model. We retrain these models to reproduce their reported results when the original model checkpoints are unavailable; however, our reproduced scores differ slightly from those originally reported.
SpaMo.
This model works by combining spatial and motion visual features using a multi-layer perceptron (MLP) and feeding them together with a prompt to an LLM for decoding and translation. Following the original setup, we set learning rate to $6\times 10^{-4}$ and train for a maximum of 100 epochs.
SEM-SLT.
SLT models have traditionally relied on glosses for training. SEM-SLT avoids this reliance by supervising on sentence embeddings instead of glosses. The full pipeline sign2(sem+text) module, with an mBART decoder from SEM-SLT, was trained with a per-device batch size of 4 for both training and validation. The learning rate was set to $1\times 10^{-5}$. Training was carried out with all other hyperparameters left at their default Hugging Face Trainer settings.
Signformer.
With the goal of edge AI, Signformer uses a combination of novel convolution, attention, and positional encoding to achieve competitive performance compared to gloss-based models while being quite smaller. Since the original paper did not specify how video features were extracted, we instead employ the S3D model pre-trained on both WLASL for ASL word recognition (Li et al., 2020) and the Kinetics-400 dataset for human action recognition (Kay et al., 2017), following the approach by (Chen et al., 2022a). We set learning rate to 0.001 and train it for a maximum of 100 epochs.
TwoStream-SLT.
Unlike most SLT methods that extract visual embeddings from raw videos only, TwoStream-SLT leverages both key points (of hands, face, and upper body) and visual embeddings from raw videos to train a novel gloss-based network for sign language recognition and SLT. We used the pretrained Sign2Gloss and the Gloss2Text modules from the TwoStream-SLT repository to reproduce their results and outputs.
Method Model G-Eval Adequacy G-Eval Fluency GEMBA
(Orig / Para) (Orig / Para)     (Orig / Para)
Signformer gpt-4.1-nano 2.54 / 2.57 3.34 / 3.48 38.45 / 38.33
qwen3:8b 2.70 / 2.71 3.49 / 3.60 74.82 / 75.27
qwq:32b 2.76 / 2.83 3.73 / 3.87 64.56 / 67.25
llama3.1:8b 3.50 / 3.56 4.65 / 4.60 57.44 / 59.58
SEM-SLT gpt-4.1-nano 2.76 / 2.91 3.26 / 3.60 47.53 / 53.17
qwen3:8b 2.97 / 3.21 3.50 / 3.95 79.79 / 81.93
qwq:32b 3.00 / 3.20 3.59 / 3.92 63.50 / 71.04
llama3.1:8b 3.42 / 3.86 4.25 / 4.63 60.63 / 65.99
SpaMo gpt-4.1-nano 3.00 / 3.06 3.66 / 3.73 55.36 / 57.31
qwen3:8b 3.33 / 3.40 4.02 / 4.12 82.26 / 83.11
qwq:32b 3.29 / 3.31 3.99 / 4.00 73.66 / 75.52
llama3.1:8b 3.95 / 4.07 4.79 / 4.72 68.32 / 70.25
TwoStream-SLT gpt-4.1-nano 3.28 / 3.33 3.77 / 3.82 64.63 / 66.69
qwen3:8b 3.67 / 3.76 4.26 / 4.38 85.67 / 85.98
qwq:32b 3.61 / 3.61 4.09 / 4.06 77.85 / 79.32
llama3.1:8b 4.20 / 4.33 4.81 / 4.70 73.85 / 75.33
Signformer (Avg) - 2.88 / 2.92 3.80 / 3.89 58.82 / 60.11
SEM-SLT (Avg) - 3.04 / 3.30 3.65 / 4.03 62.86 / 68.03
SpaMo (Avg) - 3.39 / 3.46 4.12 / 4.14 69.90 / 71.55
TwoStream-SLT (Avg) - 3.69 / 3.76 4.23 / 4.24 75.50 / 76.83
Table 2: G-Eval adequacy and fluency scores, along with GEMBA, for original and paraphrased translations on the Phoenix-2014T test set.
4.  Evaluation Results
4.1.  Impact of Paraphrasing on SLT Evaluation
One of the main limitations of lexical overlap metrics is their sensitivity to paraphrasing. We provide a comparison of the scores with and without paraphrasing for every metric and model in Table 1. Most prominently, we observe that lexical overlap metrics (BLEU, chrF, and ROUGE) are highly sensitive to variations in word order, even though our paraphrasing prompt preserves the overall sentence structure. Notably, the BLEU score drops to slightly less than half of its original value, while metrics such as chrF and ROUGE also decrease but to a much lesser extent. Interestingly, embedding-based BLEURT scores increase across all methods when comparing paraphrased translations to the original ones. This increase is particularly pronounced for SEM-SLT, increasing from $0.484$ to $0.537$. This is most likely because the SEM-SLT model is trained with a contrastive learning method using sentence-level embeddings, which allows it to generalize better to semantically equivalent paraphrases. To alleviate the sensitivity of lexical overlap metrics to paraphrasing and failure of capturing the overall semantics, we compare the metrics in a multi-reference setting, as shown in Table 1. We observe that including a multi-reference setting ensures that evaluation metrics fairly reward semantic equivalence when evaluating models that exhibit high surface-level variation, such as those subject to paraphrasing.
We present the results of LLM-based evaluators, G-Eval (Adequacy and Fluency) and GEMBA, comparing paraphrased and original translations on the Phoenix-2014T test set in Table 2. Most notably, paraphrasing generally improves G-Eval Adequacy and Fluency and GEMBA scores, consistent with the pattern observed for BLEURT. Llama3.1:8b is the only model that rates paraphrased translations as less fluent than the original ones. Among all evaluated LLMs, Llama3.1:8b tends to assign the highest Adequacy and Fluency scores, whereas gpt-4.1-nano assigns the lowest. We observe that LLM-based evaluators tend to prefer paraphrased translations over the original ones. This tendency likely arises because the paraphrases are themselves generated by an LLM, which may introduce stylistic patterns or linguistic preferences that align with the evaluator’s own training distribution, thereby biasing its judgments. Interestingly, this observation aligns with previous studies suggesting that LLMs exhibit a bias toward LLM-generated or paraphrased text (Xu et al., 2024; Liu et al., 2023). To be able to make more general statements about evaluation metrics, we show the Pearson correlation between the lexical metrics, BLEURT, and GEMBA, computed using the average scores from Table 2 in Figure 1. We notice a high correlation among lexical overlap metrics, and GEMBA is also highly correlated with BLEURT, as both incorporate learned language model representations in their scoring. Our previous observations from Tables 1 and 2 are confirmed, with lexical metrics (BLEU, chrF, ROUGE) yielding consistent rankings, while LLM-based evaluations align closely with BLEURT, ranking TwoStream-SLT highest, followed by SpaMo, SEM-SLT, and Signformer.
Figure 1: Pearson correlation between lexical metrics (BLEU, chrF, ROUGE), BLEURT, and GEMBA.
Figure 2: Examination of the sensitivity of evaluation metrics, including BLEU, chrF, ROUGE, GEMBA, and BLEURT, to hallucinations in SLT outputs on the Phoenix-2014T test set.
4.2.  Impact of Hallucination on SLT Evaluation
We first analyze how hallucinations impact automatic metrics and then examine how robust SLT models are in the case of hallucination. Figure 2 illustrates the impact of hallucinations, categorized by their severity ranking, on various evaluation metrics across four SLT models. We omitted G-Eval in this setting because its results are reported on a Likert scale.  [^8] Typically ranging from 1 (strongly disagree/poor) to 5 (strongly agree/excellent) (Likert, 1932). Moreover, we report hallucination detection results for GEMBA using gpt-4.1-nano. Notably, results show that there is a monotonic decline across all metrics, meaning that as hallucination severity level increases, translation quality drops. However, BLEU is an extremely sensitive metric; in cases such as Partial Hallucinations or Full Hallucinations, even minor rephrasings that remain semantically correct can cause a sharp decrease in BLEU scores. ROUGE and chrF decline more smoothly, since they account for partial overlaps at the subword or character level, respectively. GEMBA and BLEURT follow a similar trend, as they can still capture some semantic meaning even in extreme Full Hallucination cases where the surface wording differs substantially. However, we suspect that these metrics may under-penalize subtle hallucinations when the generated text is fluent, thereby overestimating the overall score. Similar observations regarding BLEURT’s overestimation and the data leakage issues in LLMs have been discussed by Zeng et al. (2024).
Figure 3: Evaluation of four SLT models (SEM-SLT, Signformer, SpaMo, TwoStream-SLT) across sentence length bins (1–6, 7–12, 13–18, 19–24, 25–31) on the Phoenix-2014T test set.
Different from the main focus of the paper that concentrates on evaluating evaluation metrics, here we briefly take the opportunity to investigate the behavior of SLT models when hallucinations occurs. Overall, TwoStream-SLT and SpaMo demonstrate a consistent trend across all evaluation metrics, showing less hallucinations compared to other models. In contrast, Signformer exhibits notably lower average performance, aligning with the results presented in Table 1. To analyze the extent to which SLT models produce hallucinations, we present in Table 3 a breakdown of the number of hallucinated translations according to severity level. The results reveal that gloss-based TwoStream-SLT yields the lowest Full Hallucinations (24), indicating higher robustness to severe errors, whereas Signformer produces the highest number (177), confirming our earlier observations regarding its overall vulnerability to hallucination. This finding is in line with previous work showing that gloss-free models exhibit higher hallucination rates than gloss-based models (Hamidullah et al., 2025a).
Model No Small Partial Full
Signformer 87 193 185 177
SEM-SLT 187 222 174 59
SpaMo 204 231 138 69
TwoStream-SLT 296 228 94 24
Table 3: Number of hallucinations based on severity level (No, Small, Partial, and Full Hallucinations) across SLT models.
4.3.  Impact of Sentence Length on SLT Evaluation
To highlight the limitations of text-based automatic metrics in SLT, we evaluate their performance across five sentence-length bins, as shown in Figure 3. We omitted G-Eval (Fluency) for this experiment as we are concerned with the overall translation quality. Across all evaluation metrics, performance consistently decreases with sentence length, indicating that longer sentences pose greater challenges for current SLT models, though the rate and shape of decline vary by metric. Notably, we observe that SEM-SLT ranks third according to BLEURT or LLM-based (GEMBA and G-Eval) metrics but first or second under lexical overlap ones, revealing a metric inconsistency.
Additionally, we analyze how translation quality varies with sentence length across different SLT models. Notably, gloss-based TwoStream-SLT demonstrates the strongest overall performance and exhibits the least sensitivity to longer sequences. SEM-SLT performs competitively on short sentences (1–6 tokens) but its performance declines more sharply for longer sentences, particularly on BLEU, GEMBA, and G-Eval. SpaMo shows moderate yet consistent performance, with relatively smaller drops for longer sequences compared to the other models. Finally, Signformer underperforms across all metrics and sentence-length bins, consistent with the trends observed in our earlier experiments.
5.  Qualitative Analysis
To better understand the impact of paraphrasing and hallucination on translation quality, we analyze specific examples exhibiting extreme cases of hallucination. Specifically, Partial Hallucinations and Full Hallucinations, where the GEMBA score exceeds 50. Notably, we did not find any instances of Full Hallucinations with a GEMBA score above 50 across all SLT models. Table 4 presents one illustrative example per SLT model. We observe that paraphrasing tends to increase the GEMBA score in cases of Partial Hallucinations across all models. By examining the paraphrased examples, we find that paraphrased translations are generally more fluent and natural than the original outputs, even though hallucinations persist. However, this is not always the case, as in most examples, paraphrasing either leaves the GEMBA score unchanged or leads to a decrease.
Model SpaMo
GEMBA 65 $\rightarrow$ 85
Ref. am wochenende beruhigt sich dann das wetter langsam .
(Over the weekend, the weather gradually calms down.)
Pred. am wochenende erwartet uns dann verbreitet ruhiges recht trockenes winterwetter .
Paraph. am wochenende erwartet uns weitgehend ruhiges und recht trockenes winterwetter .
Model SEM-SLT
GEMBA 65 $\rightarrow$ 85
Ref. im norden und nordosten bleibt es meist bedeckt mitunter fällt dort etwas regen .
(In the north and northeast, it mostly stays cloudy, with occasional light rain there.)
Pred. im norden und nordosten fällt regen sonst ist es meist trocken .
Paraph. im norden und nordosten regnet es, während es andernorts meist trocken bleibt .
Model TwoStream-SLT
GEMBA 75 $\rightarrow$ 85
Ref. in der südhälfte muss dazu mit nachtfrost gerechnet werden .
(In the southern half, one must expect nighttime frost.)
Pred. im süden gibt es heute nacht teilweise strengen frost .
Paraph. im süden wird es heute nacht teilweise sehr frostig sein .
Model Signformer
GEMBA 75 $\rightarrow$ 85
Ref. südlich des mains morgen verbreitet freundlich .
(South of the Main, it will be generally sunny tomorrow.)
Pred. vor allem in der südwesthälfte ist es länger freundlich .
Paraph. insbesondere in der südwestlichen hälfte bleibt es länger freundlich .
Table 4: Qualitative examples in the presence of Partial Hallucinations with paraphrased predictions.
6.  Conclusion
In this work, we present a systematic analysis of text-based automatic evaluation metrics for SLT, including traditional lexical overlap metrics such as BLEU and ROUGE, and recent embedding- and LLM-based metrics such as BLEURT, GEMBA, and G-Eval. Our results highlight the limitations of both categories: lexical metrics (BLEU, chrF, ROUGE) often fail to capture semantic adequacy, particularly under paraphrasing, whereas embedding-based BLEURT and LLM-based evaluators (G-Eval and GEMBA), despite being more semantically aware, tend to overestimate the quality of fluent but hallucinated outputs. Based on these findings, we recommend using BLEURT or LLM-based evaluators such as GEMBA and G-Eval along with traditional lexical overlap metrics. Furthermore, we advocate for the development and adoption of multimodal evaluation frameworks that extend beyond text-based metrics to provide a more holistic and comprehensive assessment of SLT outputs.
Limitations
In this work, we investigate the limitations of text-based automatic evaluation metrics in SLT through empirical experiments. While our study covers large and controlled settings, it has several limitations. First, our evaluation relies solely on automatic metrics, as no human evaluation was conducted to serve as a gold standard. This limitation stems from the lack of available native signers for large-scale annotation. Future work should include human judgments from native sign language users to establish a more reliable evaluation benchmark. Second, our experiments are restricted to the Phoenix-2014T dataset, which contains German Sign Language (DGS) data primarily focused on weather forecasts. This narrow domain may limit the generalizability of our findings to broader SLT contexts. Finally, although our results consistently highlight the weaknesses of lexical overlap metrics, we rely primarily on BLEU for quantitative evaluation. However, BLEU scores can be unreliable when computed on small test sets: prior work has shown that BLEU’s correlation with human judgments becomes unstable when sample sizes are low (Mathur et al., 2020). To mitigate this issue, we applied bootstrap resampling to estimate confidence intervals, which provides a more robust assessment of metric stability and significance. We emphasize that these experiments are part of an ongoing research effort, and the findings reported here are preliminary and should not be deployed or regarded as final without community approval.
7.  Bibliographical References
[…]
  • Benkirane et al. (2024) Kenza Benkirane, Laura Gongas, Shahar Pelles, Naomi Fuchs, Joshua Darmon, Pontus Stenetorp, David Ifeoluwa Adelani, and Eduardo Sánchez. 2024. Machine translation hallucination detection for low and high resource languages using large language models. In Findings of the Association for Computational Linguistics: EMNLP 2024, pages 9647–9665, Miami, Florida, USA. Association for Computational Linguistics.
  • Camgoz et al. (2018) Necati Cihan Camgoz, Simon Hadfield, Oscar Koller, Hermann Ney, and Richard Bowden. 2018. Neural sign language translation. In 2018 IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 7784–7793.
  • Camgoz et al. (2020) Necati Cihan Camgoz, Oscar Koller, Simon Hadfield, and Richard Bowden. 2020. Sign language transformers: Joint end-to-end sign language recognition and translation. In IEEE Conference on Computer Vision and Pattern Recognition (CVPR).
  • Chen et al. (2022a) Yutong Chen, Fangyun Wei, Xiao Sun, Zhirong Wu, and Stephen Lin. 2022a. A simple multi-modality transfer learning baseline for sign language translation. 2022 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pages 5110–5120.
  • Chen et al. (2022b) Yutong Chen, Ronglai Zuo, Fangyun Wei, Yu Wu, Shujie Liu, and Brian Mak. 2022b. Two-stream network for sign language recognition and translation. In Proceedings of the 36th International Conference on Neural Information Processing Systems, NIPS ’22, Red Hook, NY, USA. Curran Associates Inc.
  • Chen et al. (2024) Zhigang Chen, Benjia Zhou, Jun Li, Jun Wan, Zhen Lei, Ning Jiang, Quan Lu, and Guoqing Zhao. 2024. Factorized learning assisted with large language model for gloss-free sign language translation. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), pages 7071–7081, Torino, Italia. ELRA and ICCL.
  • Cheng et al. (2023) Yiting Cheng, Fangyun Wei, Bao Jianmin, Dong Chen, and Wen Qiang Zhang. 2023. Cico: Domain-aware sign language retrieval via cross-lingual contrastive learning. In CVPR.
  • et al. (2024) OpenAI et al. 2024. Gpt-4o system card. ArXiv.
  • Fomicheva et al. (2022) Marina Fomicheva, Shuo Sun, Erick Fonseca, Chrysoula Zerva, Frédéric Blain, Vishrav Chaudhary, Francisco Guzmán, Nina Lopatina, Lucia Specia, and André F. T. Martins. 2022. MLQE-PE: A multilingual quality estimation and post-editing dataset. In Proceedings of the Thirteenth Language Resources and Evaluation Conference, pages 4963–4974, Marseille, France. European Language Resources Association.
  • Freitag et al. (2024) Markus Freitag, Nitika Mathur, Daniel Deutsch, Chi-Kiu Lo, Eleftherios Avramidis, Ricardo Rei, Brian Thompson, Frederic Blain, Tom Kocmi, Jiayi Wang, David Ifeoluwa Adelani, Marianna Buchicchio, Chrysoula Zerva, and Alon Lavie. 2024. Are LLMs breaking MT metrics? results of the WMT24 metrics shared task. In Proceedings of the Ninth Conference on Machine Translation, pages 47–81, Miami, Florida, USA. Association for Computational Linguistics.
  • Gong et al. (2024) Jia Gong, Lin Geng Foo, Yixuan He, Hossein Rahmani, and Jun Liu. 2024. Llms are good sign language translators. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pages 18362–18372.
  • Guan et al. (2025) Mo Guan, Yan Wang, Guangkun Ma, Jiarui Liu, and Mingzu Sun. 2025. Mska: Multi-stream keypoint attention network for sign language recognition and translation. Pattern Recogn., 165(C).
  • Guerreiro et al. (2023) Nuno M. Guerreiro, Duarte M. Alves, Jonas Waldendorf, Barry Haddow, Alexandra Birch, Pierre Colombo, and André F. T. Martins. 2023. Hallucinations in large multilingual translation models. Transactions of the Association for Computational Linguistics, 11:1500–1517.
  • Hamidullah et al. (2025a) Yasser Hamidullah, Koel Dutta Chowdury, Yusser Al-Ghussin, Shakib Yazdani, Cennet Oguz, Josef van Genabith, and Cristina España-Bonet. 2025a. Grounding or guessing? visual signals for detecting hallucinations in sign language translation. ArXiv.
  • Hamidullah et al. (2024) Yasser Hamidullah, Josef van Genabith, and Cristina España-Bonet. 2024. Sign language translation with sentence embedding supervision. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 2: Short Papers), pages 425–434, Bangkok, Thailand. Association for Computational Linguistics.
  • Hamidullah et al. (2025b) Yasser Hamidullah, Shakib Yazdani, Cennet Oguz, Josef van Genabith, and Cristina España-Bonet. 2025b. Sonar-slt: Multilingual sign language translation via language-agnostic sentence embedding supervision. In Proceedings of the Tenth Conference on Machine Translation (WMT 2025), pages 301–313, Suzhou, China. Association for Computational Linguistics.
  • Hwang et al. (2025) Eui Jun Hwang, Sukmin Cho, Junmyeong Lee, and Jong C. Park. 2025. An efficient gloss-free sign language translation using spatial configurations and motion dynamics with LLMs. In Proceedings of the 2025 Conference of the Nations of the Americas Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), pages 3901–3920, Albuquerque, New Mexico. Association for Computational Linguistics.
  • Imai et al. (2025) Saki Imai, Mert Inan, Anthony B. Sicilia, and Malihe Alikhani. 2025. Silverscore: Semantically-aware embeddings for sign language generation evaluation. In Proceedings of the 15th International Conference on Recent Advances in Natural Language Processing - Natural Language Processing in the Generative AI era, pages 452–461, Varna, Bulgaria. INCOMA Ltd., Shoumen, Bulgaria.
  • Ji et al. (2023) Ziwei Ji, Nayeon Lee, Rita Frieske, Tiezheng Yu, Dan Su, Yan Xu, Etsuko Ishii, Ye Jin Bang, Andrea Madotto, and Pascale Fung. 2023. Survey of hallucination in natural language generation. ACM Comput. Surv., 55(12).
  • Kay et al. (2017) Will Kay, João Carreira, Karen Simonyan, Brian Zhang, Chloe Hillier, Sudheendra Vijayanarasimhan, Fabio Viola, Tim Green, Trevor Back, Apostol Natsev, Mustafa Suleyman, and Andrew Zisserman. 2017. The kinetics human action video dataset. ArXiv, abs/1705.06950.
  • Kim et al. (2024) Jung-Ho Kim, Mathew Huerta-Enochian, Changyong Ko, and Du Hui Lee. 2024. SignBLEU: Automatic evaluation of multi-channel sign language translation. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), pages 14796–14811, Torino, Italia. ELRA and ICCL.
  • Kocmi and Federmann (2023) Tom Kocmi and Christian Federmann. 2023. GEMBA-MQM: Detecting translation quality error spans with GPT-4. In Proceedings of the Eighth Conference on Machine Translation, pages 768–775, Singapore. Association for Computational Linguistics.
  • Li et al. (2020) Dongxu Li, Cristian Rodriguez, Xin Yu, and Hongdong Li. 2020. Word-level deep sign language recognition from video: A new large-scale dataset and methods comparison. In The IEEE Winter Conference on Applications of Computer Vision, pages 1459–1469.
  • Likert (1932) Rensis Likert. 1932. A technique for the measurement of attitudes. Archives of Psychology, 22(140):1–55.
  • Lin (2004) Chin-Yew Lin. 2004. ROUGE: A package for automatic evaluation of summaries. In Text Summarization Branches Out, pages 74–81, Barcelona, Spain. Association for Computational Linguistics.
  • Liu et al. (2023) Yang Liu, Dan Iter, Yichong Xu, Shuohang Wang, Ruochen Xu, and Chenguang Zhu. 2023. G-eval: NLG evaluation using gpt-4 with better human alignment. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, pages 2511–2522, Singapore. Association for Computational Linguistics.
  • Mathur et al. (2020) Nitika Mathur, Timothy Baldwin, and Trevor Cohn. 2020. Tangled up in BLEU: Reevaluating the evaluation of automatic machine translation evaluation metrics. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 4984–4997, Online. Association for Computational Linguistics.
  • Müller et al. (2023) Mathias Müller, Malihe Alikhani, Eleftherios Avramidis, Richard Bowden, Annelies Braffort, Necati Cihan Camgöz, Sarah Ebling, Cristina España-Bonet, Anne Göhring, Roman Grundkiewicz, Mert Inan, Zifan Jiang, Oscar Koller, Amit Moryossef, Annette Rios, Dimitar Shterionov, Sandra Sidler-Miserez, Katja Tissi, and Davy Van Landuyt. 2023. Findings of the second WMT shared task on sign language translation (WMT-SLT23). In Proceedings of the Eighth Conference on Machine Translation, pages 68–94, Singapore. Association for Computational Linguistics.
  • Müller et al. (2022) Mathias Müller, Sarah Ebling, Eleftherios Avramidis, Alessia Battisti, Michèle Berger, Richard Bowden, Annelies Braffort, Necati Cihan Camgöz, Cristina España-Bonet, Roman Grundkiewicz, Zifan Jiang, Oscar Koller, Amit Moryossef, Regula Perrollaz, Sabine Reinhard, Annette Rios, Dimitar Shterionov, Sandra Sidler-Miserez, Katja Tissi, and Davy Van Landuyt. 2022. Findings of the First WMT Shared Task on Sign Language Translation (WMT-SLT22). In Proceedings of the Seventh Conference on Machine Translation, pages 744–772, Abu Dhabi. Association for Computational Linguistics.
  • Papineni et al. (2002) Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002. Bleu: a method for automatic evaluation of machine translation. In Proceedings of the 40th Annual Meeting of the Association for Computational Linguistics, pages 311–318, Philadelphia, Pennsylvania, USA. Association for Computational Linguistics.
  • Popović (2015) Maja Popović. 2015. chrF: character n-gram F-score for automatic MT evaluation. In Proceedings of the Tenth Workshop on Statistical Machine Translation, pages 392–395. Association for Computational Linguistics.
  • Post (2018) Matt Post. 2018. A call for clarity in reporting BLEU scores. In Proceedings of the Third Conference on Machine Translation: Research Papers, pages 186–191, Brussels, Belgium. Association for Computational Linguistics.
  • Sato et al. (2024) Ayako Sato, Kyotaro Nakajima, Hwichan Kim, Zhousi Chen, and Mamoru Komachi. 2024. TMU-HIT’s submission for the WMT24 quality estimation shared task: Is GPT-4 a good evaluator for machine translation? In Proceedings of the Ninth Conference on Machine Translation, pages 529–534, Miami, Florida, USA. Association for Computational Linguistics.
  • Sellam et al. (2020) Thibault Sellam, Dipanjan Das, and Ankur Parikh. 2020. BLEURT: Learning robust metrics for text generation. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 7881–7892, Online. Association for Computational Linguistics.
  • Stokoe (1980) William C. Stokoe. 1980. Sign language structure. Annual Review of Anthropology, 9:365–390.
  • Tan et al. (2025) Sihan Tan, Taro Miyazaki, and Kazuhiro Nakadai. 2025. Multilingual gloss-free sign language translation: Towards building a sign language foundation model. In Proceedings of the 63rd Annual Meeting of the Association for Computational Linguistics (Volume 2: Short Papers), pages 553–561, Vienna, Austria. Association for Computational Linguistics.
  • Tong et al. (2025) Tony Cheng Tong, Sirui He, Zhiwen Shao, and Dit-Yan Yeung. 2025. G-veval: A versatile metric for evaluating image and video captions using gpt-4o. Proceedings of the AAAI Conference on Artificial Intelligence, 39(7):7419–7427.
  • Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Brian Ichter, Fei Xia, Ed H. Chi, Quoc V. Le, and Denny Zhou. 2022. Chain-of-thought prompting elicits reasoning in large language models. In Proceedings of the 36th International Conference on Neural Information Processing Systems, NIPS ’22, Red Hook, NY, USA. Curran Associates Inc.
  • Wong et al. (2024) Ryan Wong, Necati Cihan Camgoz, and Richard Bowden. 2024. Sign2GPT: Leveraging large language models for gloss-free sign language translation. In The Twelfth International Conference on Learning Representations.
  • Xu et al. (2024) Wenda Xu, Guanglei Zhu, Xuandong Zhao, Liangming Pan, Lei Li, and William Wang. 2024. Pride and prejudice: LLM amplifies self-bias in self-refinement. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 15474–15492, Bangkok, Thailand. Association for Computational Linguistics.
  • Yang (2024) Eta Yang. 2024. Signformer is all you need: Towards edge ai for sign language. ArXiv.
  • Yazdani et al. (2025) Shakib Yazdani, Josef Van Genabith, and Cristina España-Bonet. 2025. Continual learning in multilingual sign language translation. In Proceedings of the 2025 Conference of the Nations of the Americas Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), pages 10923–10938, Albuquerque, New Mexico. Association for Computational Linguistics.
  • Yin et al. (2022) Aoxiong Yin, Zhou Zhao, Weike Jin, Meng Zhang, Xingshan Zeng, and Xiaofei He. 2022. Mlslt: Towards multilingual sign language translation. In 2022 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pages 5099–5109.
  • Yin et al. (2023) Aoxiong Yin, Tianyun Zhong, Li Tang, Weike Jin, Tao Jin, and Zhou Zhao. 2023. Gloss attention for gloss-free sign language translation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pages 2551–2562.
  • Zeng et al. (2024) Xianfeng Zeng, Yijin Liu, Fandong Meng, and Jie Zhou. 2024. Towards multiple references era – addressing data leakage and limited reference diversity in machine translation evaluation. In Findings of the Association for Computational Linguistics: ACL 2024, pages 11939–11951, Bangkok, Thailand. Association for Computational Linguistics.
  • Zhang et al. (2023) Biao Zhang, Mathias Müller, and Rico Sennrich. 2023. SLTUNET: A simple unified model for sign language translation. In The Eleventh International Conference on Learning Representations.
  • Zhou et al. (2021) Hao Zhou, Wengang Zhou, Weizhen Qi, Junfu Pu, and Houqiang Li. 2021. Improving sign language translation with monolingual data by sign back-translation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pages 1316–1325.
Appendix A Prompts
A.1.  Hallucination Detection
System:You are an expert evaluator of Sign Language Translation (SLT) outputs.You will be given:- A human reference translation- A model predicted translationYour task is to determine **how severe** the hallucination is in thepredicted translation.Definition:A word in the translated text is considered a hallucination if it introducesinformation that is completely unrelated to the source text.Assign one severity level according to these guidelines:• No hallucination: The translated text does not contain any hallucinated words.• Small hallucination: The translated text contains 1-2 hallucinated words.• Partial hallucination: The translated text includes at least 3hallucinated words, but not all words are hallucinated.• Full hallucination: Nearly all words in the translated text arehallucinated, with the exception of perhaps 1-2 words.Note: The labels are mutually exclusive; for example, a translation witha partial hallucination does not qualify as a full hallucination.User:Reference Translation: {ref_text}Predicted Translation: {mt_text}Provide exactly one of the following labels as your response. Do not include anyadditional text or explanation:• No hallucination• Small hallucination• Partial hallucination• Full hallucination
A.2.  G-Eval
You will be given a generated translation for a short sign language video segment,along with reference translation.Your task is to rate the generated translation based on its Adequacy and Fluencyin capturing the intended meaning of the original signing as conveyedin the reference translation.Evaluation Criteria:Score Range: 1 to 5 (integer) - The generated translation should:- Correctly convey the meaning expressed in the reference translation.- Include all key information without introducing unrelated or incorrect content.- Be written in clear and natural language.- Stay true to the meaning of the original signing.Evaluation Dimensions:1. Adequacy — Does the translation correctly convey the meaning?2. Fluency — Is the translation grammatically correct, coherent, and easy to read?Evaluation Steps:1. Examine the reference translation to understand themeaning, key elements, and tone.2. Read the generated translation thoroughly.3. Compare the generated translation with the references and judge how well it: - Captures meaning accurately. - Covers all important elements. - Maintains grammatical fluency. - Stays faithful to the original intent.4. Penalize for: - Missing key elements. - Introducing unrelated or incorrect details. - Awkward or unclear phrasing. - Changing the tone or meaning.5. Assign an integer score from 1 to 5 for each dimension: - Adequacy score - Fluency scoreReference Translation:{{Reference}}Generated Translation:{{Translation}}Format of Output:You should first give an explanation for each score, then end withtwo separate sentences:..... The Adequacy score: {{Adequacy_score}}..... The Fluency score: {{fluency_score}}
A.3.  GEMBA
Score the following translation from German Sign Language to Germanwith respect to the human reference, on a continuous scale from 0 to 100,where 0 means no meaning preserved and 100 means perfect meaning and grammar.Scoring guidelines:0–10: Incorrect translation (no relation to the source meaning).11–29: Contains a few correct keywords, but the overall meaningdiffers significantly.30–50: Major mistakes that distort meaning.51–69: Understandable and conveys the main meaning but includes noticeablegrammatical or lexical errors.70–90: Closely preserves the semantics of the source sentence,with only minor issues.91–100: Perfect translation—fully accurate, fluent, and natural.Human reference:{{Reference}}Model translation:{{Translation}}..... Score: {{score}}

### v2

@@
 Sign languages are the primary communication systems for millions of deaf and hard-of-hearing people across the globe.
-They are highly expressive visual languages that convey meaning through hand signs, facial expressions, mouthings, and body posture (Stokoe, 1980).
-Automatic sign language translation (SLT) aims to translate sign language videos into spoken language text to help bridge the communication gap between the deaf and hearing communities (Camgoz et al., 2018).
+They are highly expressive visual languages that convey meaning through hand signs, facial expressions, mouthings, and body posture (stokoe1980sign).
+Automatic sign language translation (SLT) aims to translate sign language videos into spoken language text to help bridge the communication gap between the deaf and hearing communities (Camgoz_2018_CVPR).
 However, evaluating SLT systems remains a significant challenge, since current evaluation practices rely solely on textual outputs, offering little insight into how effectively the visual-linguistic nature of sign language is captured.
@@
 As a result, current SLT models typically rely on text-based automatic evaluation metrics.
-Consequently, a valid measure of translation quality that accounts for this cross-modal grounding is still missing in SLT (Müller et al., 2023).
-In addition to the limitations of current evaluation metrics, SLT systems are also prone to hallucinations (Zhang et al., 2023; Hamidullah et al., 2025a), fluent but incorrect translations that misrepresent the signed input.
+Consequently, a valid measure of translation quality that accounts for this cross-modal grounding is still missing in SLT (muller-etal-2023-findings).
+In addition to the limitations of current evaluation metrics, SLT systems are also prone to hallucinations (zhang2023sltunet; hamidullah2025grounding), fluent but incorrect translations that misrepresent the signed input.
 Evaluating how metrics respond to such cases is crucial for understanding their reliability.
-When translating from sign language to spoken language text, traditional lexical overlap machine translation (MT) metrics such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) are commonly adopted.
+When translating from sign language to spoken language text, traditional lexical overlap machine translation (MT) metrics such as BLEU (papineni-etal-2002-bleu) and ROUGE (lin-2004-rouge) are commonly adopted.
 However, these text-based metrics fall short in capturing the multimodal complexity of sign language.
 To address this gap, recent studies have proposed evaluation metrics tailored to sign language tasks.
-SignBLEU (Kim et al., 2024), for instance, operates at the gloss level and aligns better with human judgments, while SiLVERScore (Imai et al., 2025) introduces an embedding-based approach for assessing sign language generation.
+SignBLEU (kim-etal-2024-signbleu), for instance, operates at the gloss level and aligns better with human judgments, while SiLVERScore (imai-EtAl:2025:RANLP2) introduces an embedding-based approach for assessing sign language generation.
 Yet, SignBLEU depends on gloss annotations, and SiLVERScore has not been evaluated in the context of SLT.
-Beyond these, embedding-based and large language model (LLM)-based metrics developed for text-to-text MT—such as BLEURT (Sellam et al., 2020) and GEMBA (Kocmi and Federmann, 2023) have shown strong correlations with human judgments in text-to-text MT (Freitag et al., 2024).
+Beyond these, embedding-based and large language model (LLM)-based metrics developed for text-to-text MT—such as BLEURT (sellam-etal-2020-bleurt) and GEMBA (kocmi-federmann-2023-gemba) have shown strong correlations with human judgments in text-to-text MT (freitag-etal-2024-llms).
 Nevertheless, these approaches remain underexplored in SLT, and there is still no comprehensive analysis of their strengths and limitations for evaluating SLT systems.
@@
 SLT models have traditionally relied on lexical overlap MT metrics for automatic translation.
-Specifically, BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) have been widely used following the neural SLT work by Camgoz et al.
-(2018).
-Following Müller et al.
-(2022), studies in SLT have adopted the embedding-based metric BLEURT (Sellam et al., 2020).
+Specifically, BLEU (papineni-etal-2002-bleu) and ROUGE (lin-2004-rouge) have been widely used following the neural SLT work by Camgoz_2018_CVPR.
+Following muller-etal-2022-findings, studies in SLT have adopted the embedding-based metric BLEURT (sellam-etal-2020-bleurt).
 However, these metrics do not take the visual source input into account.
-Recently, Kim et al.
-(2024) proposed SignBLEU, an evaluation metric designed to reduce information loss when assessing SLT outputs.
+Recently, kim-etal-2024-signbleu proposed SignBLEU, an evaluation metric designed to reduce information loss when assessing SLT outputs.
 By operating at the gloss level, it achieves stronger alignment with human judgments than prior text-based metrics.
 Nevertheless, SignBLEU remains constrained to the text modality and relies on manually curated gloss annotations, which limits scalability.
-More recently, Imai et al.
-(2025) introduced SiLVERScore, an embedding-based metric that evaluates sign language generation in a multimodal semantic space.
+More recently, imai-EtAl:2025:RANLP2 introduced SiLVERScore, an embedding-based metric that evaluates sign language generation in a multimodal semantic space.
 While it shows promising results, SiLVERScore was not evaluated on SLT.
 The progress and advancements in LLMs have motivated research into their potential applications for automated assessment across various tasks.
-Liu et al.
-(2023) pioneered LLM-based automated assessment by proposing G-Eval, a framework that uses Chain-of-Thought reasoning (Wei et al., 2022) along with task descriptions and evaluation metrics to evaluate performance on tasks including text summarization and dialogue generation.
-Following this direction, Kocmi and Federmann (2023) proposed GEMBA, a GPT-based evaluation metric to assess the MT quality both with- and without reference in a zero-shot prompting fashion.
-Recently, inspired by G-Eval, Tong et al.
-(2025) extended the framework to image and video caption evaluation using GPT-4o, achieving a high correlation with human judgments.
+liu-etal-2023-g pioneered LLM-based automated assessment by proposing G-Eval, a framework that uses Chain-of-Thought reasoning (10.5555/3600270.3602070) along with task descriptions and evaluation metrics to evaluate performance on tasks including text summarization and dialogue generation.
+Following this direction, kocmi-federmann-2023-gemba proposed GEMBA, a GPT-based evaluation metric to assess the MT quality both with- and without reference in a zero-shot prompting fashion.
+Recently, inspired by G-Eval, Tong_He_Shao_Yeung_2025 extended the framework to image and video caption evaluation using GPT-4o, achieving a high correlation with human judgments.
 $2.$  Sign Language Translation
@@
 SLT models are often categorized into gloss-free and methods with gloss supervision.
-Camgoz et al.
-(2018) pioneered SLT by framing it as a neural machine translation (NMT) task, and further extending it by leveraging transformers for end-to-end SLT (Camgoz et al., 2020).
-MSKA-SLT (Guan et al., 2025) sets a new state-of-the-art on the Phoenix-2014T (Camgoz et al., 2018) and CSL-Daily (Zhou et al., 2021) benchmarks as a gloss-based model, with BLEU scores of $29.03$ and $25.52$, respectively.
+Camgoz_2018_CVPR pioneered SLT by framing it as a neural machine translation (NMT) task, and further extending it by leveraging transformers for end-to-end SLT (camgoz2020sign).
+MSKA-SLT (10.1016/j.patcog.2025.111602) sets a new state-of-the-art on the Phoenix-2014T (Camgoz_2018_CVPR) and CSL-Daily (Zhou_2021_CVPR) benchmarks as a gloss-based model, with BLEU scores of $29.03$ and $25.52$, respectively.
 The approach models keypoints extracted from the face, body, and both hands using an attention mechanism, and is jointly optimized with a CTC loss and a self-distillation loss.
@@
 In contrast, recent approaches attempt to model SLT at the sentence level and a gloss-free fashion.
-Building on this objective, recent works have incorporated LLMs into SLT frameworks (Gong et al., 2024; Wong et al., 2024; Chen et al., 2024), employed contrastive learning objectives (Cheng et al., 2023), replaced gloss supervision with sentence-level embeddings (Hamidullah et al., 2024), and extended SLT to multilingual scenarios (Yin et al., 2022; Yazdani et al., 2025; Tan et al., 2025; Hamidullah et al., 2025b).
+Building on this objective, recent works have incorporated LLMs into SLT frameworks (gong2024signllm; wong2024sign2gpt; chen-etal-2024-factorized), employed contrastive learning objectives (cheng2023cico), replaced gloss supervision with sentence-level embeddings (hamidullah2024slt), and extended SLT to multilingual scenarios (yin2022multislt; yazdani-etal-2025-continual; tan-etal-2025-multilingual; hamidullah-etal-2025-sonar).
 2.3.  Hallucination
-A model is said to hallucinate when it generates content that either lacks logical coherence or diverges from the source material, asserting information or events that are unfounded or inconsistent with the known facts (Ji et al., 2023).
-While hallucination detection has received increasing attention in the MT community, with studies on multilingual comparisons (Guerreiro et al., 2023) and LLM-based approaches for both low- and high-resource languages (Benkirane et al., 2024), research on hallucination in SLT remains limited.
-Zhang et al.
-(2023) reported, based on manual analysis, that SLT models such as SLTUNET suffer greatly from hallucinations, where the generated translations often show limited correlation with the sign video.
-More recently, Hamidullah et al.
-(2025a) proposed a hallucination detection framework and further demonstrated that gloss-free SLT models are particularly prone to severe hallucinations compared to their gloss-based counterparts.
+A model is said to hallucinate when it generates content that either lacks logical coherence or diverges from the source material, asserting information or events that are unfounded or inconsistent with the known facts (survey-of-hallucination).
+While hallucination detection has received increasing attention in the MT community, with studies on multilingual comparisons (guerreiro-etal-2023-hallucinations) and LLM-based approaches for both low- and high-resource languages (benkirane-etal-2024-machine), research on hallucination in SLT remains limited.
+zhang2023sltunet reported, based on manual analysis, that SLT models such as SLTUNET suffer greatly from hallucinations, where the generated translations often show limited correlation with the sign video.
+More recently, hamidullah2025grounding proposed a hallucination detection framework and further demonstrated that gloss-free SLT models are particularly prone to severe hallucinations compared to their gloss-based counterparts.
 These findings highlight the need for a deeper understanding of how evaluation metrics respond to hallucinated outputs and how robust current SLT models are under such conditions.
@@
 We use the Phoenix-2014T  [^1] Phoenix-2014T was selected because it is the most widely used benchmark in SLT research and contains gloss-level annotations necessary for evaluating gloss-based models.
-dataset (Camgoz et al., 2018), covering weather forecasts in German sign language (DGS) to evaluate and compare the performance of various models in the above aspects.
+dataset (Camgoz_2018_CVPR), covering weather forecasts in German sign language (DGS) to evaluate and compare the performance of various models in the above aspects.
 Impact of paraphrasing on SLT evaluation.
@@
 Yet, there has been no systematic investigation into how such hallucinations affect translation quality or the reliability of automatic SLT evaluation metrics.
-Motivated by findings from Benkirane et al.
-(2024), which demonstrate the effectiveness of LLMs for hallucination detection in MT, we employ Llama-3-70B to identify hallucinated outputs in a reference-based setting.
+Motivated by findings from benkirane-etal-2024-machine, which demonstrate the effectiveness of LLMs for hallucination detection in MT, we employ Llama-3-70B to identify hallucinated outputs in a reference-based setting.
 [^2] The prompt used for hallucination detection is available in Appendix A.1.
@@
 3.2.  Evaluation Metrics
-For our analysis, we consider the metrics as per Müller et al.
-(2022, 2023), specifically BLEU [^3] BLEU|nrefs:1|bs:25|tok:none|eff:no| case:mixed|smooth:exp|version:2.5.1 (via SacreBLEU; Post, 2018) for lexical overlap, ROUGE (Lin, 2004) [^4] ROUGE|metrics:rougeL|nrefs:1| stemmer:true|bs:25|version:0.1.2 for recall-oriented n-gram overlap, chrF (Popović, 2015)555chrF|nrefs:1|bs:25|case:mixed|eff:yes| nc:6|nw:2|space:no|version:2.5.1 that uses character $n$-grams, and embedding-based BLEURT (Sellam et al., 2020) [^6] BLEURT v0.0.2 using checkpoint BLEURT-20.
+For our analysis, we consider the metrics as per muller-etal-2022-findings; muller-etal-2023-findings, specifically BLEU [^3] BLEU|nrefs:1|bs:25|tok:none|eff:no| case:mixed|smooth:exp|version:2.5.1 (via SacreBLEU; post-2018-call) for lexical overlap, ROUGE (lin-2004-rouge) [^4] ROUGE|metrics:rougeL|nrefs:1| stemmer:true|bs:25|version:0.1.2 for recall-oriented n-gram overlap, chrF (popovic-2015-chrf)555chrF|nrefs:1|bs:25|case:mixed|eff:yes| nc:6|nw:2|space:no|version:2.5.1 that uses character $n$-grams, and embedding-based BLEURT (sellam-etal-2020-bleurt) [^6] BLEURT v0.0.2 using checkpoint BLEURT-20.
 for semantic quality.
 We use bootstrap resampling for statistical significance reporting.
-We also extend the traditional range of SLT evaluation metrics by presenting a sign language adapted version of G-Eval (Liu et al., 2023).
+We also extend the traditional range of SLT evaluation metrics by presenting a sign language adapted version of G-Eval (liu-etal-2023-g).
 G-Eval introduces a framework that leverages prompting techniques to generate evaluation scores that strongly correlate with human preferences.
-The prompt is organized into three modules: (1) evaluation dimensions, which specify the aspects to be judged; (2) step-by-step reasoning, where Chain-of-Thought (CoT) (Wei et al., 2022) guides the LLM through the evaluation procedure; and (3) scoring with references, which constrains the output format and incorporates human translations as ground truth.
-Following Sato et al.
-(2024), we prompt the LLMs to evaluate the translation quality of the models in Section 3.3 for Adequacy and Fluency, each rated on a 5-point Likert scale (1155).
-Following Fomicheva et al.
-(2022); Kocmi and Federmann (2023), we further include GEMBA zero-shot Direct Assessment (Kocmi and Federmann, 2023), where the LLM is prompted to evaluate each translation hypothesis and provide a quality score ranging from 0 (completely incorrect) to $100$ (perfect translation).
+The prompt is organized into three modules: (1) evaluation dimensions, which specify the aspects to be judged; (2) step-by-step reasoning, where Chain-of-Thought (CoT) (10.5555/3600270.3602070) guides the LLM through the evaluation procedure; and (3) scoring with references, which constrains the output format and incorporates human translations as ground truth.
+Following sato-etal-2024-tmu, we prompt the LLMs to evaluate the translation quality of the models in Section 3.3 for Adequacy and Fluency, each rated on a 5-point Likert scale (1155).
+Following fomicheva-etal-2022-mlqe; kocmi-federmann-2023-gemba, we further include GEMBA zero-shot Direct Assessment (kocmi-federmann-2023-gemba), where the LLM is prompted to evaluate each translation hypothesis and provide a quality score ranging from 0 (completely incorrect) to $100$ (perfect translation).
 Our LLM experimental setup includes four models: GPT-4.1-nano, Qwen3-8B, Llama-3.1-8B, and QWQ-32B.
@@
 SLT models are either gloss-based or gloss-free.
-We include three gloss-free models: SpaMo (Hwang et al., 2025), SEM-SLT (Hamidullah et al., 2024), and Signformer (Yang, 2024).
-We also include TwoStream-SLT (Chen et al., 2022b) as a gloss-based model.
+We include three gloss-free models: SpaMo (hwang-etal-2025-spamo), SEM-SLT (hamidullah2024slt), and Signformer (yang2024signformerneededgeai).
+We also include TwoStream-SLT (10.5555/3600270.3601510) as a gloss-based model.
 We retrain these models to reproduce their reported results when the original model checkpoints are unavailable; however, our reproduced scores differ slightly from those originally reported.
@@
 With the goal of edge AI, Signformer uses a combination of novel convolution, attention, and positional encoding to achieve competitive performance compared to gloss-based models while being quite smaller.
-Since the original paper did not specify how video features were extracted, we instead employ the S3D model pre-trained on both WLASL for ASL word recognition (Li et al., 2020) and the Kinetics-400 dataset for human action recognition (Kay et al., 2017), following the approach by (Chen et al., 2022a).
+Since the original paper did not specify how video features were extracted, we instead employ the S3D model pre-trained on both WLASL for ASL word recognition (li2020word) and the Kinetics-400 dataset for human action recognition (kay2017kineticshumanactionvideo), following the approach by (Chen2022ASM).
 We set learning rate to 0.001 and train it for a maximum of 100 epochs.
@@
 This tendency likely arises because the paraphrases are themselves generated by an LLM, which may introduce stylistic patterns or linguistic preferences that align with the evaluator’s own training distribution, thereby biasing its judgments.
-Interestingly, this observation aligns with previous studies suggesting that LLMs exhibit a bias toward LLM-generated or paraphrased text (Xu et al., 2024; Liu et al., 2023).
+Interestingly, this observation aligns with previous studies suggesting that LLMs exhibit a bias toward LLM-generated or paraphrased text (xu-etal-2024-pride; liu-etal-2023-g).
 To be able to make more general statements about evaluation metrics, we show the Pearson correlation between the lexical metrics, BLEURT, and GEMBA, computed using the average scores from Table 2 in Figure 1.
@@
 We omitted G-Eval in this setting because its results are reported on a Likert scale.
-[^8] Typically ranging from 1 (strongly disagree/poor) to 5 (strongly agree/excellent) (Likert, 1932).
+[^8] Typically ranging from 1 (strongly disagree/poor) to 5 (strongly agree/excellent) (likert1932technique).
 Moreover, we report hallucination detection results for GEMBA using gpt-4.1-nano.
@@
 However, we suspect that these metrics may under-penalize subtle hallucinations when the generated text is fluent, thereby overestimating the overall score.
-Similar observations regarding BLEURT’s overestimation and the data leakage issues in LLMs have been discussed by Zeng et al.
-(2024).
+Similar observations regarding BLEURT’s overestimation and the data leakage issues in LLMs have been discussed by zeng-etal-2024-towards.
 Figure 3: Evaluation of four SLT models (SEM-SLT, Signformer, SpaMo, TwoStream-SLT) across sentence length bins (1–6, 7–12, 13–18, 19–24, 25–31) on the Phoenix-2014T test set.
@@
 The results reveal that gloss-based TwoStream-SLT yields the lowest Full Hallucinations (24), indicating higher robustness to severe errors, whereas Signformer produces the highest number (177), confirming our earlier observations regarding its overall vulnerability to hallucination.
-This finding is in line with previous work showing that gloss-free models exhibit higher hallucination rates than gloss-based models (Hamidullah et al., 2025a).
+This finding is in line with previous work showing that gloss-free models exhibit higher hallucination rates than gloss-based models (hamidullah2025grounding).
 Model No Small Partial Full
@@
 Finally, although our results consistently highlight the weaknesses of lexical overlap metrics, we rely primarily on BLEU for quantitative evaluation.
-However, BLEU scores can be unreliable when computed on small test sets: prior work has shown that BLEU’s correlation with human judgments becomes unstable when sample sizes are low (Mathur et al., 2020).
+However, BLEU scores can be unreliable when computed on small test sets: prior work has shown that BLEU’s correlation with human judgments becomes unstable when sample sizes are low (mathur-etal-2020-tangled).
 To mitigate this issue, we applied bootstrap resampling to estimate confidence intervals, which provides a more robust assessment of metric stability and significance.
@@
 7.  Bibliographical References
-[…]
-  • Benkirane et al.
-(2024) Kenza Benkirane, Laura Gongas, Shahar Pelles, Naomi Fuchs, Joshua Darmon, Pontus Stenetorp, David Ifeoluwa Adelani, and Eduardo Sánchez.
-2024.
-Machine translation hallucination detection for low and high resource languages using large language models.
-In Findings of the Association for Computational Linguistics: EMNLP 2024, pages 9647–9665, Miami, Florida, USA.
-Association for Computational Linguistics.
-  • Camgoz et al.
-(2018) Necati Cihan Camgoz, Simon Hadfield, Oscar Koller, Hermann Ney, and Richard Bowden.
-2018.
-Neural sign language translation.
-In 2018 IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 7784–7793.
-  • Camgoz et al.
-(2020) Necati Cihan Camgoz, Oscar Koller, Simon Hadfield, and Richard Bowden.
-2020.
-Sign language transformers: Joint end-to-end sign language recognition and translation.
-In IEEE Conference on Computer Vision and Pattern Recognition (CVPR).
-  • Chen et al.
-(2022a) Yutong Chen, Fangyun Wei, Xiao Sun, Zhirong Wu, and Stephen Lin.
-2022a.
-A simple multi-modality transfer learning baseline for sign language translation.
-2022 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pages 5110–5120.
-  • Chen et al.
-(2022b) Yutong Chen, Ronglai Zuo, Fangyun Wei, Yu Wu, Shujie Liu, and Brian Mak.
-2022b.
-Two-stream network for sign language recognition and translation.
-In Proceedings of the 36th International Conference on Neural Information Processing Systems, NIPS ’22, Red Hook, NY, USA.
-Curran Associates Inc.
-  • Chen et al.
-(2024) Zhigang Chen, Benjia Zhou, Jun Li, Jun Wan, Zhen Lei, Ning Jiang, Quan Lu, and Guoqing Zhao.
-2024.
-Factorized learning assisted with large language model for gloss-free sign language translation.
-In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), pages 7071–7081, Torino, Italia.
-ELRA and ICCL.
-  • Cheng et al. (2023) Yiting Cheng, Fangyun Wei, Bao Jianmin, Dong Chen, and Wen Qiang Zhang. 2023. Cico: Domain-aware sign language retrieval via cross-lingual contrastive learning. In CVPR.
-  • et al. (2024) OpenAI et al. 2024. Gpt-4o system card. ArXiv.
-  • Fomicheva et al.
-(2022) Marina Fomicheva, Shuo Sun, Erick Fonseca, Chrysoula Zerva, Frédéric Blain, Vishrav Chaudhary, Francisco Guzmán, Nina Lopatina, Lucia Specia, and André F.
-T.
-Martins.
-2022.
-MLQE-PE: A multilingual quality estimation and post-editing dataset.
-In Proceedings of the Thirteenth Language Resources and Evaluation Conference, pages 4963–4974, Marseille, France.
-European Language Resources Association.
-  • Freitag et al.
-(2024) Markus Freitag, Nitika Mathur, Daniel Deutsch, Chi-Kiu Lo, Eleftherios Avramidis, Ricardo Rei, Brian Thompson, Frederic Blain, Tom Kocmi, Jiayi Wang, David Ifeoluwa Adelani, Marianna Buchicchio, Chrysoula Zerva, and Alon Lavie.
-2024.
-Are LLMs breaking MT metrics?
-results of the WMT24 metrics shared task.
-In Proceedings of the Ninth Conference on Machine Translation, pages 47–81, Miami, Florida, USA.
-Association for Computational Linguistics.
-  • Gong et al.
-(2024) Jia Gong, Lin Geng Foo, Yixuan He, Hossein Rahmani, and Jun Liu.
-2024.
-Llms are good sign language translators.
-In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pages 18362–18372.
-  • Guan et al.
-(2025) Mo Guan, Yan Wang, Guangkun Ma, Jiarui Liu, and Mingzu Sun.
-2025.
-Mska: Multi-stream keypoint attention network for sign language recognition and translation.
-Pattern Recogn., 165(C).
-  • Guerreiro et al.
-(2023) Nuno M.
-Guerreiro, Duarte M.
-Alves, Jonas Waldendorf, Barry Haddow, Alexandra Birch, Pierre Colombo, and André F.
-T.
-Martins.
-2023.
-Hallucinations in large multilingual translation models.
-Transactions of the Association for Computational Linguistics, 11:1500–1517.
-  • Hamidullah et al.
-(2025a) Yasser Hamidullah, Koel Dutta Chowdury, Yusser Al-Ghussin, Shakib Yazdani, Cennet Oguz, Josef van Genabith, and Cristina España-Bonet.
-2025a.
-Grounding or guessing?
-visual signals for detecting hallucinations in sign language translation.
-ArXiv.
-  • Hamidullah et al.
-(2024) Yasser Hamidullah, Josef van Genabith, and Cristina España-Bonet.
-2024.
-Sign language translation with sentence embedding supervision.
-In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 2: Short Papers), pages 425–434, Bangkok, Thailand.
-Association for Computational Linguistics.
-  • Hamidullah et al.
-(2025b) Yasser Hamidullah, Shakib Yazdani, Cennet Oguz, Josef van Genabith, and Cristina España-Bonet.
-2025b.
-Sonar-slt: Multilingual sign language translation via language-agnostic sentence embedding supervision.
-In Proceedings of the Tenth Conference on Machine Translation (WMT 2025), pages 301–313, Suzhou, China.
-Association for Computational Linguistics.
-  • Hwang et al.
-(2025) Eui Jun Hwang, Sukmin Cho, Junmyeong Lee, and Jong C.
-Park.
-2025.
-An efficient gloss-free sign language translation using spatial configurations and motion dynamics with LLMs.
-In Proceedings of the 2025 Conference of the Nations of the Americas Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), pages 3901–3920, Albuquerque, New Mexico.
-Association for Computational Linguistics.
-  • Imai et al.
-(2025) Saki Imai, Mert Inan, Anthony B.
-Sicilia, and Malihe Alikhani.
-2025.
-Silverscore: Semantically-aware embeddings for sign language generation evaluation.
-In Proceedings of the 15th International Conference on Recent Advances in Natural Language Processing - Natural Language Processing in the Generative AI era, pages 452–461, Varna, Bulgaria.
-INCOMA Ltd., Shoumen, Bulgaria.
-  • Ji et al.
-(2023) Ziwei Ji, Nayeon Lee, Rita Frieske, Tiezheng Yu, Dan Su, Yan Xu, Etsuko Ishii, Ye Jin Bang, Andrea Madotto, and Pascale Fung.
-2023.
-Survey of hallucination in natural language generation.
-ACM Comput.
-Surv., 55(12).
-  • Kay et al.
-(2017) Will Kay, João Carreira, Karen Simonyan, Brian Zhang, Chloe Hillier, Sudheendra Vijayanarasimhan, Fabio Viola, Tim Green, Trevor Back, Apostol Natsev, Mustafa Suleyman, and Andrew Zisserman.
-2017.
-The kinetics human action video dataset.
-ArXiv, abs/1705.06950.
-  • Kim et al.
-(2024) Jung-Ho Kim, Mathew Huerta-Enochian, Changyong Ko, and Du Hui Lee.
-2024.
-SignBLEU: Automatic evaluation of multi-channel sign language translation.
-In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), pages 14796–14811, Torino, Italia.
-ELRA and ICCL.
-  • Kocmi and Federmann (2023) Tom Kocmi and Christian Federmann.
-2023.
-GEMBA-MQM: Detecting translation quality error spans with GPT-4.
-In Proceedings of the Eighth Conference on Machine Translation, pages 768–775, Singapore.
-Association for Computational Linguistics.
-  • Li et al.
-(2020) Dongxu Li, Cristian Rodriguez, Xin Yu, and Hongdong Li.
-2020.
-Word-level deep sign language recognition from video: A new large-scale dataset and methods comparison.
-In The IEEE Winter Conference on Applications of Computer Vision, pages 1459–1469.
-  • Likert (1932) Rensis Likert. 1932. A technique for the measurement of attitudes. Archives of Psychology, 22(140):1–55.
-  • Lin (2004) Chin-Yew Lin. 2004. ROUGE: A package for automatic evaluation of summaries. In Text Summarization Branches Out, pages 74–81, Barcelona, Spain. Association for Computational Linguistics.
-  • Liu et al.
-(2023) Yang Liu, Dan Iter, Yichong Xu, Shuohang Wang, Ruochen Xu, and Chenguang Zhu.
-2023.
-G-eval: NLG evaluation using gpt-4 with better human alignment.
-In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, pages 2511–2522, Singapore.
-Association for Computational Linguistics.
-  • Mathur et al.
-(2020) Nitika Mathur, Timothy Baldwin, and Trevor Cohn.
-2020.
-Tangled up in BLEU: Reevaluating the evaluation of automatic machine translation evaluation metrics.
-In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 4984–4997, Online.
-Association for Computational Linguistics.
-  • Müller et al.
-(2023) Mathias Müller, Malihe Alikhani, Eleftherios Avramidis, Richard Bowden, Annelies Braffort, Necati Cihan Camgöz, Sarah Ebling, Cristina España-Bonet, Anne Göhring, Roman Grundkiewicz, Mert Inan, Zifan Jiang, Oscar Koller, Amit Moryossef, Annette Rios, Dimitar Shterionov, Sandra Sidler-Miserez, Katja Tissi, and Davy Van Landuyt.
-2023.
-Findings of the second WMT shared task on sign language translation (WMT-SLT23).
-In Proceedings of the Eighth Conference on Machine Translation, pages 68–94, Singapore.
-Association for Computational Linguistics.
-  • Müller et al.
-(2022) Mathias Müller, Sarah Ebling, Eleftherios Avramidis, Alessia Battisti, Michèle Berger, Richard Bowden, Annelies Braffort, Necati Cihan Camgöz, Cristina España-Bonet, Roman Grundkiewicz, Zifan Jiang, Oscar Koller, Amit Moryossef, Regula Perrollaz, Sabine Reinhard, Annette Rios, Dimitar Shterionov, Sandra Sidler-Miserez, Katja Tissi, and Davy Van Landuyt.
-2022.
-Findings of the First WMT Shared Task on Sign Language Translation (WMT-SLT22).
-In Proceedings of the Seventh Conference on Machine Translation, pages 744–772, Abu Dhabi.
-Association for Computational Linguistics.
-  • Papineni et al.
-(2002) Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu.
-2002.
-Bleu: a method for automatic evaluation of machine translation.
-In Proceedings of the 40th Annual Meeting of the Association for Computational Linguistics, pages 311–318, Philadelphia, Pennsylvania, USA.
-Association for Computational Linguistics.
-  • Popović (2015) Maja Popović.
-2015.
-chrF: character n-gram F-score for automatic MT evaluation.
-In Proceedings of the Tenth Workshop on Statistical Machine Translation, pages 392–395.
-Association for Computational Linguistics.
-  • Post (2018) Matt Post.
-2018.
-A call for clarity in reporting BLEU scores.
-In Proceedings of the Third Conference on Machine Translation: Research Papers, pages 186–191, Brussels, Belgium.
-Association for Computational Linguistics.
-  • Sato et al.
-(2024) Ayako Sato, Kyotaro Nakajima, Hwichan Kim, Zhousi Chen, and Mamoru Komachi.
-2024.
-TMU-HIT’s submission for the WMT24 quality estimation shared task: Is GPT-4 a good evaluator for machine translation?
-In Proceedings of the Ninth Conference on Machine Translation, pages 529–534, Miami, Florida, USA.
-Association for Computational Linguistics.
-  • Sellam et al.
-(2020) Thibault Sellam, Dipanjan Das, and Ankur Parikh.
-2020.
-BLEURT: Learning robust metrics for text generation.
-In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 7881–7892, Online.
-Association for Computational Linguistics.
-  • Stokoe (1980) William C. Stokoe. 1980. Sign language structure. Annual Review of Anthropology, 9:365–390.
-  • Tan et al.
-(2025) Sihan Tan, Taro Miyazaki, and Kazuhiro Nakadai.
-2025.
-Multilingual gloss-free sign language translation: Towards building a sign language foundation model.
-In Proceedings of the 63rd Annual Meeting of the Association for Computational Linguistics (Volume 2: Short Papers), pages 553–561, Vienna, Austria.
-Association for Computational Linguistics.
-  • Tong et al.
-(2025) Tony Cheng Tong, Sirui He, Zhiwen Shao, and Dit-Yan Yeung.
-2025.
-G-veval: A versatile metric for evaluating image and video captions using gpt-4o.
-Proceedings of the AAAI Conference on Artificial Intelligence, 39(7):7419–7427.
-  • Wei et al.
-(2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Brian Ichter, Fei Xia, Ed H.
-Chi, Quoc V.
-Le, and Denny Zhou.
-2022.
-Chain-of-thought prompting elicits reasoning in large language models.
-In Proceedings of the 36th International Conference on Neural Information Processing Systems, NIPS ’22, Red Hook, NY, USA.
-Curran Associates Inc.
-  • Wong et al.
-(2024) Ryan Wong, Necati Cihan Camgoz, and Richard Bowden.
-2024.
-Sign2GPT: Leveraging large language models for gloss-free sign language translation.
-In The Twelfth International Conference on Learning Representations.
-  • Xu et al.
-(2024) Wenda Xu, Guanglei Zhu, Xuandong Zhao, Liangming Pan, Lei Li, and William Wang.
-2024.
-Pride and prejudice: LLM amplifies self-bias in self-refinement.
-In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 15474–15492, Bangkok, Thailand.
-Association for Computational Linguistics.
-  • Yang (2024) Eta Yang. 2024. Signformer is all you need: Towards edge ai for sign language. ArXiv.
-  • Yazdani et al.
-(2025) Shakib Yazdani, Josef Van Genabith, and Cristina España-Bonet.
-2025.
-Continual learning in multilingual sign language translation.
-In Proceedings of the 2025 Conference of the Nations of the Americas Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), pages 10923–10938, Albuquerque, New Mexico.
-Association for Computational Linguistics.
-  • Yin et al.
-(2022) Aoxiong Yin, Zhou Zhao, Weike Jin, Meng Zhang, Xingshan Zeng, and Xiaofei He.
-2022.
-Mlslt: Towards multilingual sign language translation.
-In 2022 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pages 5099–5109.
-  • Yin et al.
-(2023) Aoxiong Yin, Tianyun Zhong, Li Tang, Weike Jin, Tao Jin, and Zhou Zhao.
-2023.
-Gloss attention for gloss-free sign language translation.
-In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pages 2551–2562.
-  • Zeng et al.
-(2024) Xianfeng Zeng, Yijin Liu, Fandong Meng, and Jie Zhou.
-2024.
-Towards multiple references era – addressing data leakage and limited reference diversity in machine translation evaluation.
-In Findings of the Association for Computational Linguistics: ACL 2024, pages 11939–11951, Bangkok, Thailand.
-Association for Computational Linguistics.
-  • Zhang et al.
-(2023) Biao Zhang, Mathias Müller, and Rico Sennrich.
-2023.
-SLTUNET: A simple unified model for sign language translation.
-In The Eleventh International Conference on Learning Representations.
-  • Zhou et al.
-(2021) Hao Zhou, Wengang Zhou, Weizhen Qi, Junfu Pu, and Houqiang Li.
-2021.
-Improving sign language translation with monolingual data by sign back-translation.
-In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pages 1316–1325.
 Appendix A Prompts